OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y =3/2 x + 4 and passes through the point (−6, 3). y =3/2 x − 1 y =3/2 x + 3 y = −2/3x − 1 y = −2/3x + 3

OpenStudy (anonymous):

y =3/2 x + 4 Perpendicular lines have negative reciprocal slope. So first you need to find the slope of the line that I wrote above and then find the negative reciprocal of it.

OpenStudy (anonymous):

After that you use point slope form to find the equation you want \[y-y_a=m(x-x_a)\\m=slope\\(x_a,y_a)=(−6, 3)\]